# StreakNet-Arch: An Anti-scattering Network-based Architecture for Underwater Carrier LiDAR-Radar Imaging

### v1

@@ I Introduction
    StreakNet-Arch: Enabling Real-Time Imaging. We introduce StreakNet-Arch, a novel end-to-end binary classification architecture that revolutionizes the UCLR’s signal processing. This approach empowers the UCLR with real-time imaging capabilities for the first time.
@@
    DBC-Attention: Optimized for Underwater Environments. We enhance the UCLR’s signal processing with Self-Attention networks. Further, we propose DBC-Attention, a groundbreaking variant specifically optimized for underwater imaging tasks. Experimental results conclusively demonstrate DBC-Attention’s superiority over the standard Self-Attention approach.
@@
    Learned Bandpass Filtering with Streak-Tube Camera Integration. We propose a method to embed streak-tube camera images directly into the attention network. This embedded representation effectively functions as a learned bandpass filter, as demonstrated by our experiments.
@@
    Dataset. We released a large-scale dataset containing 2,695,168 real-world underwater 3D point cloud data captured by streak-tube camera, which facilitates further development of Underwater laser imaging signal processing techniques.
@@ II-B Attention Mechanism
In the past decade, the attention mechanism has played an increasingly important role in computer vision and natural language processing. In 2014, Mnih V. et al. [43] pioneered the use of attention mechanism into neural networks, predicting crucial regions through policy gradient recursion and updating the entire network end-to-end. Later, various works [44, 45] adopted a similar strategy for attention in vision, with recurrent neural networks (RNNs) serving as necessary tools. Hu J. et al. proposed SENet [36], presenting a novel channel-attention network that implicitly and adaptively predicts potential key features.
A significant shift came in 2017 with the introduction of the Self-Attention mechanism by Vaswani et al [35]. This advancement revolutionized Natural Language Processing (NLP) [41, 42]. In 2018, Wang et al. [37] took the lead in introducing Self-Attention to computer vision. Notably, Hu et al. (2018) also proposed a novel channel-attention network (SENet) within this timeframe. Recently, various deep Self-Attention networks (Visual Transformers, ViTs) [38, 39, 40, 46, 47] have appeared, showcasing the immense potential of attention-based models.
Attention mechanisms can also be applied to the enhancement of underwater image processing [5, 48, 49]. In 2023, Peng L. et al. introduced the U-shape Transformer, pioneering the incorporation of self-attention mechanisms into underwater image enhancement [5]. They proposed a Transformer module that fuses multi-scale features across channels, and a spatial module for global feature modeling. This innovation enhances the network’s focus on areas of more severe attenuation in both color channels and spatial regions. Mehnaz U. at el. proposed an innovative Underwater window-based Transformer Generative Adversarial Network (UwTGAN) aimed at enhancing underwater image quality for computer vision applications in marine settings [48]. Pramanick A. at el. propose a framework called X-CAUNET that considered wavelength of light in underwater conditions by using cross-attention transformers [49].
@@ III-B FD Embedding Layer
$f_{s}=\frac{N_{s}}{T_{\text{full}}}.$ (1)
@@
$\begin{split}\mathbf{u^{\prime}}_{\text{echo}}&=\text{IEO}(\mathbf{u}_{\text{echo}}),\\ \mathbf{u^{\prime}}_{\text{tem}}&=\text{IEO}(\mathbf{u}_{\text{tem}}).\end{split}$ (5)
@@ III-C Attention Analysis Method
$\begin{split}{A}_{i}&=\frac{{A^{\prime}}_{i}-\min\{{A^{\prime}}_{i}\}}{\max\{{A^{\prime}}_{i}\}-\min\{{A^{\prime}}_{i}\}},\\ \mathbf{A}&=\left(A_{i}\right)_{1\times 2L},\end{split}$ (9)
@@ III-D Double Branch Cross Attention Backbone
Eq. 10-12 together form the basic block of DBC-Attention. Similar to the Transformer architecture [35], DBC-Attention can use a multi-head attention approach when calculating scores.
@@ III-E Imaging Head
The Imaging Head consists of two main data paths: the denoising path and the imaging path. The denoising path is used to determine whether the feature tensor of the input echo signal of the network contains a target, producing a mask map. This aims to replace the handcraft threshold used in traditional algorithms and can be modeled as a binary classification task. The imaging path builds upon traditional imaging methods but incorporates an equivalent filter learned through AAM during the filtering process. This learned filter replaces traditional handcrafted bandpass filters, resulting in candidate gray and distance maps. Finally, the mask map from the denoising path is multiplied element-wise with the gray and distance maps generated by the imaging path to produce the final imaging outputs.
@@
$\begin{split}\mathbf{Y}^{\prime}=\text{Concat}\left(\mathbf{Y}_{\text{echo}},\mathbf{Y}_{\text{tem}}\right),\\ \mathbf{Y}^{\top}=\text{SiLU}\left(\mathbf{W}\mathbf{Y}^{\prime\top}+b\right),\end{split}$ (13)
@@
Then, multiply $\bm{\mu}_{\text{echo}}$ by the spectrum of template signal $\mathbf{u}_{\text{tem}}$, perform frequency-domain matched filtering (Eq. 18), and transform back to the time domain using inverse fast Fourier transform (Eq. 19).
$\bm{\psi}=\bm{\mu}_{\text{echo}}\odot\mathbf{u}_{\text{tem}}.$ (18)
$\mathbf{v}_{f}=\mathcal{IFFT}(\bm{\psi},N_{\text{FFT}})[0:N_{s}],$ (19)
@@
$\begin{split}i&=\text{argmax}(\mathbf{v}_{f}),\\ t&=i\frac{1}{f_{s}}+t_{G},\\ \mathbf{CG}(j,i)&=\max(\mathbf{v}_{f}),\\ \mathbf{CD}(j,i)&=\frac{c}{n}\cdot\frac{t}{2},\end{split}$ (20)
@@
$\begin{split}\mathbf{G}&=\mathbf{CG}\odot\mathbf{M},\\ \mathbf{D}&=\mathbf{CD}\odot\mathbf{M}.\end{split}$ (21)
@@ IV-A Model training
Figure 5: Learning rate variation curve during the training phase.
Figure 6: Loss variation curve during the training phase.
@@ IV-B StreakNet-Arch exhibits superior anti-scattering capabilities compared to traditional imaging methods
$\begin{split}TP&=\sum_{i}\sum_{j}\hat{M}_{ij}\land M_{ij},\\ FP&=\sum_{i}\sum_{j}\hat{M}_{ij}\land\neg M{ij},\\ FN&=\sum_{i}\sum_{j}\neg\hat{M}_{ij}\land\neg M{ij},\\ P&=\frac{TP}{TP+FP},\\ R&=\frac{TP}{TP+FN},\\ F_{1}&=\frac{2\cdot P\cdot R}{P+R}.\end{split}$ (23)
@@
(b) Traditional 15 m
(c) Traditional 13 m
(d) Traditional 10 m
(e) StreakNet-Emb 20 m
(f) StreakNet-Emb 15 m
(g) StreakNet-Emb 13 m
(h) StreakNet-Emb 10 m
(i) StreakNetv2-Emb 20 m
(j) StreakNetv2-Emb-15m
(k) StreakNetv2-Emb 13 m
(l) StreakNetv2-Emb 10 m
(m) StreakNet 20 m
(n) StreakNet 15 m
(o) StreakNet 13 m
(p) StreakNet 10 m
(q) StreakNetv2 20 m
(r) StreakNetv2 15 m
(s) StreakNetv2 13 m
(t) StreakNetv2 10 m
@@
(b) Traditional 15 m
(c) Traditional 13 m
(d) Traditional 10 m
(e) StreakNet-Emb 20 m
(f) StreakNet-Emb 15 m
(g) StreakNet-Emb 13 m
(h) StreakNet-Emb 10 m
(i) StreakNetv2-Emb 20 m
(j) StreakNetv2-Emb 15 m
(k) StreakNetv2-Emb 13 m
(l) StreakNetv2-Emb 10 m
(m) StreakNet 20 m
(n) StreakNet 15 m
(o) StreakNet 13 m
(p) StreakNet 10 m
(q) StreakNetv2 20 m
(r) StreakNetv2 15 m
(s) StreakNetv2 13 m
(t) StreakNetv2 10 m
@@ IV-C StreakNet-Arch is more suitable for real-time imaging than Traditional Imaging Methods
$\begin{split}\text{AIT}_{\text{traditional}}&=\frac{1}{N}\sum_{i=1}^{N}t_{i1}+t_{i2}\\ &=\frac{1}{N}\sum_{i=1}^{N}it_{i1}\\ &=\frac{N+1}{2}t_{m}.\end{split}$ (24)
@@
$\begin{split}\text{AIT}_{\text{streaknet}}&=\frac{1}{N}\sum_{i=1}^{N}t_{i1}\\ &=t_{m}.\end{split}$ (25)
@@ IV-D FD Embedding Layer is an equivalent bandpass filter
Since the carrier frequency of the detection signal is 500 MHz (see Fig. 1e), traditional imaging algorithms use a handcraft bandpass filter with a range of 450 MHz - 550 MHz during filtering. If we consider the bandpass filter from the perspective of “attention distribution”, we can think of the bandpass filter as a binary attention distribution with values of 1 for frequencies in the range of 450 MHz - 550 MHz and 0 for frequencies outside this range. Similarly, the attention distribution of the FD Embedding Layer can also be regarded as a generalized bandpass filter.
@@ IV-E DBC-Attention is more suitable for underwater optical 3D imaging than Self-Attention
To validate the superiority of DBC-Attention over Self-Attention in underwater imaging tasks, we designed and conducted ablation experiments. StreakNetv2, compared to StreakNet, replaces the backbone network from Self-Attention to DBC-Attention, while keeping the other parameters unchanged. Through the experimental results (Fig. 7 and Table III), we found:
@@
    Regarding anti-scattering performance: Except for the m-model, the $F_{1}$ scores of StreakNetv2 on s, l, and x models is higher than StreakNet, indicating that the average anti-scattering performance of DBC-Attention is superior to Self-Attention.
@@
    Regarding anti-overfitting performance: The number of network parameters of s, m, l, x models increase sequentially. The $F_{1}$ scores of StreakNet and StreakNetv2 increases from s to m and decreases thereafter, indicating varying degrees of overfitting in both architectures after the m-model. Although StreakNet’s performance is significantly higher than StreakNetv2 on the m-model, it significantly decreases on the x-model. Overall, StreakNet shows large fluctuations in anti-scattering performance from s to x, while StreakNetv2 remains relatively stable, indicating that DBC-Attention has stronger anti-overfitting performance than Self-Attention in underwater imaging tasks.
We simultaneously conducted ablation experiments with traditional imaging methods. We performed Attention Analysis on the FD Embedding Layer of both StreakNet and StreakNetv2. The results were used as equivalent filters, replacing the traditional 450 MHz - 550 MHz bandpass filter for imaging, and the imaging outcomes are shown in Fig.8e-l and Fig. 9e-l. The results on the validation set are presented in Fig. 14b and Table V.

### v2

@@ I Introduction
    We introduce StreakNet-Arch, a novel end-to-end binary classification architecture that revolutionizes the UCLR’s signal processing. This approach empowers the UCLR with real-time imaging capabilities for the first time.
@@
    We enhance the UCLR’s signal processing with Self-Attention networks. Further, we propose DBC-Attention, a groundbreaking variant specifically optimized for underwater imaging tasks. Experimental results conclusively demonstrate DBC-Attention’s superiority over the standard Self-Attention approach.
@@
    We propose a method to embed streak-tube camera images directly into the attention network. This embedded representation effectively functions as a learned bandpass filter, as demonstrated by our experiments.
@@
    We released a large-scale dataset containing 2,695,168 real-world underwater 3D point cloud data captured by streak-tube camera, which facilitates further development of Underwater laser imaging signal processing techniques.
@@ II-B Attention Mechanism
In the past decade, the attention mechanism has played an increasingly important role in computer vision and natural language processing. In 2014, Mnih V. et al. [43] pioneered the use of attention mechanism into neural networks, predicting crucial regions through policy gradient recursion and updating the entire network end-to-end. Subsequent works [44, 45] in visual attention leveraged recurrent neural networks (RNNs) as essential tools. Hu J. et al. proposed SENet [36], presenting a novel channel-attention network that implicitly and adaptively predicts potential key features. A significant shift came in 2017 with the introduction of the Self-Attention mechanism by Vaswani et al [35]. This advancement revolutionized Natural Language Processing (NLP) [41, 42]. In 2018, Wang et al. [37] took the lead in introducing Self-Attention to computer vision. Notably, Hu et al. (2018) proposed a channel-attention network (SENet) within this timeframe. Recently, various Self-Attention networks (Visual Transformers, ViTs) [38, 39, 40, 46, 47] have appeared, showcasing the immense potential of attention-based models.
Attention mechanisms can also be applied to the enhancement of underwater image processing [5, 48, 49]. In 2023, Peng L. et al. introduced the U-shape Transformer, pioneering the incorporation of self-attention mechanisms into underwater image enhancement [5]. They proposed a Transformer module that fuses multi-scale features across channels, and a spatial module for global feature modeling. This innovation enhances the network’s focus on areas of more severe attenuation in both color channels and spatial regions. Mehnaz U. et al. proposed an innovative Underwater window-based Transformer Generative Adversarial Network (UwTGAN) aimed at enhancing underwater image quality for computer vision applications in marine settings [48]. Pramanick A. at el. propose a framework that considered wavelength of light in underwater conditions by using cross-attention transformers [49].
@@ III-B FD Embedding Layer
$f_{s}=\frac{N_{s}}{T_{\text{full}}},\Delta R_{f}=\frac{fs}{N_{\text{FFT}}}.$ (1)
@@
$\mathbf{u^{\prime}}_{\text{echo}}=\text{IEO}(\mathbf{u}_{\text{echo}}),\mathbf{u^{\prime}}_{\text{tem}}=\text{IEO}(\mathbf{u}_{\text{tem}}).$ (4)
@@ III-C Attention Analysis Method
${A}_{i}=\frac{{A^{\prime}}_{i}-\min\{{A^{\prime}}_{i}\}}{\max\{{A^{\prime}}_{i}\}-\min\{{A^{\prime}}_{i}\}},\mathbf{A}=\left(A_{i}\right)_{1\times 2L},$ (8)
@@ III-D Double Branch Cross Attention Backbone
Eq. 9-11 together form the basic block of DBC-Attention. Similar to the Transformer architecture [35], DBC-Attention can use a multi-head attention approach when calculating scores.
@@ III-E Imaging Head
The Imaging Head comprises two data paths: denoising and imaging. The denoising path, modeled as a binary classification task, identifies target regions within the input feature tensor using a learned mask map, replacing traditional hand-crafted thresholds. The imaging path leverages traditional methods but incorporates a learned filter (replacing handcrafted bandpass filters) obtained through AAM during filtering. This results in candidate gray and distance maps. Finally, element-wise multiplication of the denoising mask with these maps generates the final imaging outputs.
@@
$\mathbf{Y}^{\top}=\text{SiLU}\left(\mathbf{W}\cdot\text{Concat}\left(\mathbf{Y}_{\text{echo}},\mathbf{Y}_{\text{tem}}\right)^{\top}+b\right),$ (12)
@@
Then, multiply $\bm{\mu}_{\text{echo}}$ by the spectrum of template signal $\mathbf{u}_{\text{tem}}$, perform frequency-domain matched filtering, and transform back to the time domain using inverse fast Fourier transform (Eq. 17).
$\mathbf{v}_{f}=\mathcal{IFFT}(\bm{\mu}_{\text{echo}}\odot\mathbf{u}_{\text{tem}},N_{\text{FFT}})[0:N_{s}],$ (17)
@@
$\begin{split}i=\text{argmax}(\mathbf{v}_{f}),&t=i\frac{1}{f_{s}}+t_{G},\\ \mathbf{CG}(j,i)=\max(\mathbf{v}_{f}),&\mathbf{CD}(j,i)=\frac{c}{n}\cdot\frac{t}{2},\end{split}$ (18)
@@
$\mathbf{G}=\mathbf{CG}\odot\mathbf{M},\mathbf{D}=\mathbf{CD}\odot\mathbf{M}.$ (19)
@@ IV-A Model training
Figure 5: Learning rate variation curve during the training phase.
Figure 6: Loss variation curve during the training phase.
@@ IV-B StreakNet-Arch exhibits superior anti-scattering capabilities compared to traditional imaging methods
$\begin{split}P&=\frac{\sum_{i}\sum_{j}\hat{M}_{ij}\land M_{ij}}{\sum_{i}\sum_{j}\hat{M}_{ij}\land M_{ij}+\sum_{i}\sum_{j}\hat{M}_{ij}\land\neg M{ij}},\\ R&=\frac{\sum_{i}\sum_{j}\hat{M}_{ij}\land M_{ij}}{\sum_{i}\sum_{j}\hat{M}_{ij}\land M_{ij}+\sum_{i}\sum_{j}\neg\hat{M}_{ij}\land\neg M{ij}},\\ F_{1}&=\frac{2\cdot P\cdot R}{P+R}.\end{split}$ (21)
@@
(b) StreakNet 20 m
(c) StreakNetv2 20 m
(d) Traditional 15 m
(e) StreakNet 15 m
(f) StreakNetv2 15 m
(g) Traditional 13 m
(h) StreakNet 13 m
(i) StreakNetv2 13 m
(j) Traditional 10 m
(k) StreakNet 10 m
(l) StreakNetv2 10 m
@@
(b) StreakNet 20 m
(c) StreakNetv2 20 m
(d) Traditional 15 m
(e) StreakNet 15 m
(f) StreakNetv2 15 m
(g) Traditional 13 m
(h) StreakNet 13 m
(i) StreakNetv2 13 m
(j) Traditional 10 m
(k) StreakNet 10 m
(l) StreakNetv2 10 m
@@ IV-C StreakNet-Arch is more suitable for real-time imaging than Traditional Imaging Methods
$\text{AIT}_{\text{traditional}}=\frac{N+1}{2}t_{m}=\frac{N+1}{2}t_{m}.$ (22)
@@
$\text{AIT}_{\text{streaknet}}=\frac{1}{N}\sum_{i=1}^{N}t_{i1}=t_{m}.$ (23)
@@ IV-D FD Embedding Layer is an equivalent bandpass filter
Since the carrier frequency of the detection signal is 500 MHz (see Fig. 1e), traditional imaging algorithms use a handcraft bandpass filter with a range of 450 MHz - 550 MHz during filtering. If we consider the bandpass filter from the perspective of “attention distribution”, we can think of the bandpass filter as a binary attention distribution with values of 1 for frequencies in the range of 450 MHz - 550 MHz and 0 for frequencies outside this range. The FD Embedding Layer’s attention distribution offers a similar concept, functioning as a learnable generalized bandpass filter.
@@ IV-E DBC-Attention is more suitable for underwater optical 3D imaging than Self-Attention
To demonstrate the superiority of DBC-Attention over Self-Attention in underwater imaging tasks, we conducted ablation experiments by replacing the Self-Attention module in StreakNet with DBC-Attention while keeping all other parameters unchanged. Through the experimental results (Fig. 7 and Table III), we found:
@@
    Except for the m-model, the $F_{1}$ scores of StreakNetv2 on s, l, and x models is higher than StreakNet, indicating that the average anti-scattering performance of DBC-Attention is superior to Self-Attention.
@@
    The number of network parameters of s, m, l, x models increase sequentially. The $F_{1}$ scores of StreakNet and StreakNetv2 increases from s to m and decreases thereafter, indicating varying degrees of overfitting in both architectures after the m-model. Although StreakNet’s performance is significantly higher than StreakNetv2 on the m-model, it significantly decreases on the x-model. Overall, StreakNet shows large fluctuations in anti-scattering performance from s to x, while StreakNetv2 remains relatively stable, indicating that DBC-Attention has stronger anti-overfitting performance than Self-Attention in underwater imaging tasks.
We simultaneously conducted ablation experiments with traditional imaging methods. We performed Attention Analysis on the FD Embedding Layer of both StreakNet and StreakNetv2. The results were used as equivalent filters, replacing the traditional 450 MHz - 550 MHz bandpass filter for imaging. The results on the validation set are presented in Fig. 14b and Table V.